# QKFormer: Hierarchical Spiking Transformer using Q-K Attention

### v1

@@ Abstract
Spiking Transformers, which integrate Spiking Neural Networks (SNNs) with Transformer architectures, have attracted significant attention due to their potential for energy efficiency and high performance. However, existing models in this domain still suffer from suboptimal performance. We introduce several innovations to improve the performance: i) We propose a novel spike-form Q-K attention mechanism, tailored for SNNs, which efficiently models the importance of token or channel dimensions through binary vectors with linear complexity. ii) We incorporate the hierarchical structure, which significantly benefits the performance of both the brain and artificial neural networks, into spiking transformers to obtain multi-scale spiking representation. iii) We design a versatile and powerful patch embedding module with a deformed shortcut specifically for spiking transformers. Together, we develop QKFormer, a hierarchical spiking transformer based on Q-K attention with direct training. QKFormer shows significantly superior performance over existing state-of-the-art SNN models on various mainstream datasets. Notably, with comparable size to Spikformer (66.34 M, 74.81%), QKFormer (64.96 M) achieves a groundbreaking top-1 accuracy of 85.65% on ImageNet-1k, substantially outperforming Spikformer by 10.84%. To our best knowledge, this is the first time that directly training SNNs have exceeded 85% accuracy on ImageNet-1K. The code and models are publicly available at https://github.com/ zhouchenlin2096/QKFormer.
@@ 1 Introduction
Regarded as the third generation of neural networks Maass (1997), the brain-inspired Spiking Neural Networks (SNNs) are potential competitors to Artificial Neural Networks (ANNs) due to their high biological plausibility, event-driven properties and low power consumption on neuromorphic hardware Roy et al. (2019). Spiking Transformers (Transformer-based SNNs) Zhou et al. (2023c, a); Yao et al. (2023); Zhou et al. (2023b); Wang et al. (2023), which integrate spiking neural networks with transformer architecture, have attracted significant attention. This innovative combination provides great potential to break through the performance bottleneck of SNNs. However, the performance of existing spiking transformers is still limited for vision tasks on complex datasets, such as on ImageNet.
Spiking Self Attention (SSA) Zhou et al. (2023c), the core module of spiking transformers, is a novel spike-form self-attention using sparse spike-form Query, key, and Value. However, the computational complexity (especially space complexity) of SSA, scales quadratically to the number of tokens (#tokens). In addition, directly training SNNs needs to unfold the network in the time dimension. Taken together, these two factors indicate huge space consumption burden for training a transformer-based SNN. The multi-scale, hierarchical modular structure, a prominent feature in both the human brain and deep neural networks (DNNs), plays a crucial role in human cognitive functions Wang et al. (2021) and in vision tasks solved by DNNs. However, existing direct training spiking transformers are all based on the straight-through structure with a fixed small token scale Zhou et al. (2023c); Yao et al. (2023); Wang et al. (2023), lacking hierarchical spiking representation for downstream vision tasks. The high computational complexity of SSA is the main obstacle to explore hierarchical architecture with fine-grained spiking feature maps in spiking transformers.
@@
To address these issues, we propose a novel hierarchical spiking transformer using Q-K attention for direct training, named QKFormer. Q-K attention, the key module of QKFormer, could efficiently model the importance of token or channel dimensions through binary vectors with linear complexity and only adopts two spike-form components: Query ($Q$) and Key ($K$). QKFormer conducts a hierarchical spiking representation by starting from small patches and gradually merging neighboring patches in deeper spiking transformer layers, with gradually decreasing #tokens. In addition, we design a versatile and powerful patch embedding module with a deformed shortcut specifically for spiking transformers. These merits make QKFormer the most powerful SNN model, in contrast to the previous transformer-based SNNs with spiking feature maps of a single resolution and quadratic computational complexity. The main contributions of this paper are as follows:
@@
  • We develop a novel spike-form Q-K attention mechanism, tailor-made for the spatio-temporal spiking patterns of SNNs, which can easily model the importance of token or channel dimensions with binary values. The Q-K attention has linear complexity to #tokens (or #channels) and only adopts two spike-form components: Query ($Q$) and Key ($K$).
@@
  • We design a versatile and powerful Patch Embedding module with a Deformed Shortcut (PEDS), which enhances spiking information transmission thus improving the performance of spiking transformers greatly.
@@
  • We build the hierarchical spiking transformer based on the proposed Q-K attention and PEDS in a direct training way, named QKFormer. This marks the effective exploration of hierarchical spiking representation in Transformer-based SNNs.
@@
  • Extensive experiments show that the proposed model outperforms the state-of-the-art (SOTA) SNNs on several static and neuromorphic datasets. Notably, QKFormer has achieved a significant milestone, surpassing 85% top-1 accuracy on ImageNet with 4 time steps using the direct training approach for the first time.
@@ 2 Related Work
2.1 Convolution-based Spiking Neural Networks.
At present, there are mainly two ways to obtain trained SNNs. One involves converting pre-trained ANNs to SNNs Bu et al. (2021); Li et al. (2021); Hao et al. (2023), replacing the ReLU activation function in ANN with spiking neurons, resulting in comparable performance to ANNs but with high latency. Another method is to directly train SNNsWu et al. (2018), using surrogate gradientNeftci et al. (2019)Xiao et al. (2021) to address the non-differentiability of spike excitation function during backpropagation, which results in low latency but relatively poor performance with performance degradation and gradient explosion/vanishing problems. Zheng et al. (2021) proposed a threshold-dependent batch normalization (tdBN) method, which mitigated the problem of gradient explosion/vanishing, extended directly trained SNNs to 50 layers for the first time, and achieved 67.05 % accuracy with 6 time step on ImageNet. Fang et al. (2021a) proposed the Spike-Element-Wise block, which further addressed gradient explosion and gradient vanishing problems, and prolonged the directly trained SNNs beyond a depth of 100 layers. MS-ResNet Hu et al. (2021) directly trained convolution-based SNNs with pre-activation shortcut to successfully extend the depth up to 482 layers on CIFAR10 without experiencing degradation problems.
2.2 Transformer-based Spiking Neural Networks.
Transformers Vaswani et al. (2017) have become benchmark models for natural language processing tasks since it was proposed. Later, Dosovitskiy et al. developed Vision Transformer (ViT) Dosovitskiy et al. (2020), which gradually became a mainstream visual network model with its superior performance.
@@ 3 Method
Figure 1: The overview of Q-K attention, which includes two forms: Q-K token attention (QKTA) and Q-K channel attention (QKCA).
@@ 3.1 Preliminary
3.1.1 Spiking Neuron Model
Spiking neuron is the fundamental unit of SNNs, we choose Leaky Integrate-and-Fire (LIF) model as the spiking neuron in our work. The dynamics of a LIF neuron can be formulated as follows:
$H[t]=V[t-1]+\frac{1}{\tau}\left(X[t]-\left(V[t-1]-V_{\text{reset}}\right)\right),$ (1)
$S[t]=\Theta\left(H[t]-V_{th}\right),$ (2)
$V[t]=H[t](1-S[t])+V_{\text{reset }}S[t],$ (3)
where $\tau$ is the membrane time constant, and $X[t]$ is the input current at time step $t$. When the membrane potential $H[t]$ exceeds the firing threshold $V_{th}$, the spiking neuron will trigger a spike $S[t]$. $\Theta(v)$ is the Heaviside step function, which equals to 1 when $v\geq 0$ and 0 otherwise. $V[t]$ represents the membrane potential after the triggered event, which equals to $H[t]$ if no spike is generated and otherwise equals to the reset potential $V_{reset}$.
3.1.2 Vanilla Self Attention
Vanilla self-attention (VSA) Vaswani et al. (2017) in transformers has three floating-point key components: query ($Q_{\mathcal{F}}$), key ($K_{\mathcal{F}}$), value ($V_{\mathcal{F}}$) which are calculated by learnable linear matrics and input $X$. The calculation of VSA can be formulated as follows:
$Q_{\mathcal{F}},K_{\mathcal{F}},V_{\mathcal{F}}=X(W_{Q},W_{K},W_{V}),$ (4)
$\operatorname{VSA}\left(Q_{\mathcal{F}},K_{\mathcal{F}},V_{\mathcal{F}}\right)=\operatorname{Softmax}\left(\frac{Q_{\mathcal{F}}K_{\mathcal{F}}^{\mathrm{T}}}{\sqrt{d}}\right)V_{\mathcal{F}},$ (5)
where $\mathcal{F}$ denotes the floating-point form. Both floating-point matrix multiplication and softmax operation which contains exponent calculation and division operation, do not align with the properties of SNNs.
3.1.3 Spiking Self Attention
Spikformer Zhou et al. (2023c) demonstrated a novel spike-form self-attention named Spiking Self Attention (SSA), using sparse spike-form $Q,K,V$ without softmax operation and floating-point matrix multiplication. The calculation process of SSA is formulated as follows:
$\mathrm{I}=\mathrm{SN_{I}}\left(\mathrm{BN_{I}}\left(X(W_{\mathrm{I}})\right)\right),\mathrm{I}\in(Q,K,V),$ (6)
$\operatorname{SSA}^{\prime}(Q,K,V)=\mathrm{SN}\left(QK^{\mathrm{T}}V*s\right),$ (7)
$\operatorname{SSA}(Q,K,V)=\mathrm{SN}\left(\operatorname{BN}\left(\operatorname{Linear}\left(\operatorname{SSA}^{\prime}(Q,K,V)\right)\right)\right),$ (8)
where $Q,K,V\in\mathcal{R}^{T\times N\times D}$, the spike-form $Q,K,V$ are computed by learnable linear layers. $s$ is a scaling factor. $\mathrm{SN}$ means spiking neuron layer. The calculation of SSA avoids floating-point multiplication, meeting the property of SNNs.
@@ 3.2 Q-K Attention
An overview of Q-K attention is shown in Figure 1. Both VSA and SSA use three key components and have $O(N^{2}d)$ or $O(Nd^{2})$ computational complexity, while our proposed Q-K Attention which has linear complexity and only uses two spike-form components: Query ($Q$) and Key ($K$), which are produced through learnable linear matrics.
$Q=\mathrm{SN}_{Q}\left(\mathrm{BN}\left(XW_{Q}\right)\right),K=\mathrm{SN}_{K}\left(\mathrm{BN}\left(XW_{K}\right)\right),$ (9)
@@
where $X$ is the input spiking map. According to the detailed calculation mechanism of $Q,K$, Q-K Attention can be divided into Q-K Token Attention (QKTA) and Q-K Channel Attention (QKCA).
3.2.1 Q-K Token Attention
@@
We here assume $T=1$ and single head attention for mathematical description. After obtaining spike-form $Q,K\in\mathcal{R}^{T\times N\times D}$, both $Q$ and $K$ can be formed as a spike matrix $N\times D$ ($N$ is the token number, $D$ is the channel number). Q-K Token Attention process can be formulated as follows:
${A}_{t}=\mathrm{SN}(\sum_{i=0}^{D}Q_{i,j}),\quad X^{\prime}={A}_{t}\otimes K,$ (10)
@@
where $A_{t}$ is the $N*1$ token attention vector, which models the binary importance of different tokens. $A_{t}$ is a spike-form vector, which is obtained by addition operations (row summation) of Q spike matrix and a following spiking neuron. $\otimes$ is the Hadamard product between spike tensors, which is equivalent to the mask operation. We apply the spike-form token attention vector $A_{t}$ to the $K$ spike matrix through the column mask operation (token mask), to obtain the output $X^{\prime}$ of Q-K Token Attention.
3.2.2 Q-K Channel Attention
The calculation process of Q-K channel attention is similar to the previous Q-K token attention, and can be formulated as follows:
${A}_{c}=\mathrm{SN}(\sum_{j=0}^{N}Q_{i,j}),\quad X^{\prime}={A}_{c}\otimes K,$ (11)
@@ 3.2.2 Q-K Channel Attention
where $A_{c}$ is the $1*D$ channel attention vector, which models the binary importance of different channels. $A_{t}$ is a spike-form vector, which is obtained by addition operations (column summation) of Q spike matrix and a following spiking neuron. Then, the output $X^{\prime}$ of Q-K Channel Attention is obtained by the row mask operation (channel mask) between $A_{t}$ and $K$.
$X^{\prime\prime}=\mathrm{SN}\left(\mathrm{BN}\left({\mathrm{Linear}}\left(X^{\prime}\right)\right)\right).$ (12)
@@
As shown in Formula.12, a post-linear layer is also required after obtaining $X^{\prime}$ of Q-K Token or Channel Attention. In addition, the channel dimension is $D/h$ in the multi-head Q-K attention, where $h$ is the head number. In this work, the spiking neuron uses the LIF model Fang et al. (2021a). Same with Zhou et al. (2023c), time step $T$ is an independent dimension for the spiking neuron layer. In other layers, it is merged with the batch size. We exploit QKTA in our experiments by default.
3.2.3 Discussion on Q-K Attention
@@
Methods Time Space
VSA $O(N^{2}*D)$ $O(N^{2}+ND)$
SSA Zhou et al. (2023c) $O(N^{2}*D)$ $O(N^{2}+ND)$
SDSA Yao et al. (2023) $O(ND)$ $O(ND)$
QKTA (ours) $O(D)$ $O(N)$
QKCA (ours) $O(N)$ $O(D)$
Computational complexity: As shown in Table 1, the time complexity of Q-K attention varies based on the implementation approach. Specifically, when utilizing spike-form broadcasted element-wise multiplication, $\otimes$, the time complexity can reach up to $O(N*D)$. When applying mask operation, the time complexity of Q-K attention is only $O(N)$ or $O(D)$. The space complexity of Q-K attention with the whole process is $O(N*D)$ at most, which is caused by the self-storage consumption Q and K matrix. In terms of the space complexity of attention operation, Q-K attention only requires an extra $1*D$ or $N*1$ space to store the attention vector with the space complexity of $O(N)$ or $O(D)$. The linear complexity of Q-K attention makes it possible to successfully explore the large-scale hierarchical architecture SNN model. Energy efficiency analysis: All the elements in Q-K attention are in spike-form (either 0 or 1), thus linear multiplication is transformed into sparse addition. Mask operation can be implemented on neuromorphic chips through addressing algorithms Richter et al. (2022) or AND logic operationsPei et al. (2019) with negligible power consumption. Compared with SSA, Q-K attention is much more energy-efficient since Q-K attention only adopts two spike-form components for spike [0, 1] operation without the V input and scale operation of SSA and thus has fewer synaptic computing.

### v2

@@ Abstract
Spiking Transformers, which integrate Spiking Neural Networks (SNNs) with Transformer architectures, have attracted significant attention due to their potential for low energy consumption and high performance. However, there remains a substantial gap in performance between SNNs and Artificial Neural Networks (ANNs). To narrow this gap, we have developed QKFormer, a direct training spiking transformer with the following features: i) Linear complexity and high energy efficiency, the novel spike-form Q-K attention module efficiently models the token or channel attention through binary vectors and enables the construction of larger models. ii) Multi-scale spiking representation, achieved by a hierarchical structure with the different number of tokens across blocks. iii) Spiking Patch Embedding with Deformed Shortcut (SPEDS), enhances spiking information transmission and integration, thus improving overall performance. It is shown that QKFormer achieves significantly superior performance over existing state-of-the-art SNN models on various mainstream datasets. Notably, with comparable size to Spikformer (66.34 M, 74.81%), QKFormer (64.96 M) achieves a groundbreaking top-1 accuracy of 85.65% on ImageNet-1k, substantially outperforming Spikformer by 10.84%. To our best knowledge, this is the first time that directly training SNNs have exceeded 85% accuracy on ImageNet-1K. The code and models are available at https://github.com/zhouchenlin2096/QKFormer.
@@ 1 Introduction
Regarded as the third generation of neural networks [1], the brain-inspired Spiking Neural Networks (SNNs) are potential competitors to Artificial Neural Networks (ANNs) due to their high biological plausibility and high energy efficiency attributed to their event-driven properties [2]. Transformer, originally designed for natural language processing [3], has flourished in various computer vision tasks, including image classification [4, 5], object detection [6, 7, 8] and semantic segmentation [9, 10]. Spiking Transformers (Transformer-based SNNs) [11, 12, 13, 14, 15], which integrate spiking neural networks with transformer architecture, have attracted significant attention. This innovative combination provides great potential to develop advanced AI algorithms with high performance and low energy consumption.
As the architecture of the transformers is essential to the model’s performance [4, 5, 16, 8, 10], designing new architectures for transformer-based SNNs is quite challenging in terms of space requirements for the following reasons [11, 13, 15]. i). Spiking Self Attention (SSA) [11], the core module of spiking transformers, encodes Query, Key, and Value with sparse spikes. However, the computational complexity (especially space complexity) of SSA scales quadratically to the number of tokens (#tokens), and is the main obstacle to explore architecture that incorporate multi-level features. ii). SNNs process data across the time domain, necessitating a high level of computational and memory resources. This combination leads to considerable consumption of computational resources, making the training process highly demanding in terms of both memory and processing power.
@@
To address these issues, we propose QKFormer with three innovations. i) Q-K attention with linear complexity and high energy efficiency. ii) A hierarchical architecture with decreasing number of tokens across blocks. iii) A novel patch embedding with deformed shortcut module. The linear complexity of Q-K attention is originated from the binary spike-form vector attention. This design lower the energy consumption and the space requirement. The hierarchical architecture starts from small patches and gradually merges neighboring patches in deeper spiking transformer layers with gradually decreasing #tokens, which enables multi-level spiking feature representation and benefits the model performance. The patch embedding with deformed shortcut facilitates spiking information transmission and integration. These merits make QKFormer achieve state-of-the-art performance in the SNN domain, in contrast to the previous transformer-based SNNs with spiking feature maps of a single resolution. Our main contributions are as follows:
@@
1) We develop a novel spike-form Q-K attention mechanism, tailor-made for the spatio-temporal spiking patterns of SNNs, which can easily model the importance of token or channel dimensions with binary values. The Q-K attention has linear complexity to #tokens (or #channels) and only adopts two spike-form components: Query ($\mathbf{Q}$) and Key ($\mathbf{K}$).
@@
2) We design a versatile and powerful Spiking Patch Embedding with Deformed Shortcut (SPEDS) module, which enhances spiking information transmission and integration thus improving the performance of spiking transformers significantly.
@@
3) We build a direct-training hierarchical spiking transformer with different number of tokens across blocks, incorporating Q-K attention and SPEDS, named QKFormer. This marks the effective exploration of hierarchical spiking representation in Transformer-based SNNs.
@@
4) Extensive experiments show that the proposed model outperforms the state-of-the-art (SOTA) SNNs on various static and neuromorphic datasets. Notably, QKFormer has achieved a significant milestone, surpassing 85% top-1 accuracy on ImageNet with 4 time steps using the direct training approach for the first time.
Figure 1: Illustration of Q-K attention with the two versions of Q-K token attention (QKTA) and Q-K channel attention (QKCA). The inputs are binary spikes and there are only sparse additions and mask operations in Q-K attention. As a spike-driven module, Q-K attention efficiently models the token or channel attention through spike-form binary vectors, performing linear complexity to #tokens (or #channels) and high energy efficiency. Spiking Neuron (SN) in this work adopts the Leaky-Integrate-and-Fire (LIF) model, which is shown in Appendix. 6.1.
@@ 2 Related Work
Learning Methods of Spiking Neural Networks. At present, there are mainly two ways to obtain trained SNNs. One involves converting pre-trained ANNs to SNNs (ANN2SNN) [17, 18, 19, 20, 21, 22, 23], replacing the ReLU activation function in ANN with spiking neurons. However, This converted SNN suffers from long converting time steps and constraints on the original ANN design. Another method is to directly train SNNs[24], using surrogate gradient[25, 26, 27, 28] to address the non-differentiability of spike excitation function during backpropagation. The direct training method has received more attention due to its low latency and supporting flexible architectural exploration.
@@ 3.1 Preliminary
Vanilla Self Attention. Vanilla self-attention (VSA) [3] in transformers has three floating-point key components: query ($\mathbf{Q}_{\mathcal{F}}$), key ($\mathbf{K}_{\mathcal{F}}$), value ($\mathbf{V}_{\mathcal{F}}$) which are calculated by learnable linear matrics and input $\mathbf{X}$. The calculation of VSA can be formulated as follows:
$\mathbf{Q}_{\mathcal{F}},\mathbf{K}_{\mathcal{F}},\mathbf{V}_{\mathcal{F}}=\mathbf{X}(\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V}),$ (1)
$\operatorname{VSA}\left(\mathbf{Q}_{\mathcal{F}},\mathbf{K}_{\mathcal{F}},\mathbf{V}_{\mathcal{F}}\right)=\operatorname{Softmax}\left(\frac{\mathbf{Q}_{\mathcal{F}}\mathbf{K}_{\mathcal{F}}^{\mathrm{T}}}{\sqrt{d}}\right)\mathbf{V}_{\mathcal{F}},$ (2)
where $\mathcal{F}$ denotes the floating-point form. Both floating-point matrix multiplication and softmax operation which contains exponent calculation and division, do not align with the properties of SNNs.
Spiking Self Attention. Spikformer [11] demonstrated a novel spike-form self-attention named Spiking Self Attention (SSA), using sparse spike-form $\mathbf{Q,K,V}$ without softmax operation and floating-point matrix multiplication. The calculation process of SSA is formulated as follows:
$\mathbf{I}=\mathrm{SN}_{I}\left(\mathrm{BN}_{I}\left(\mathbf{X}(\mathbf{W}_{I})\right)\right),\mathbf{I}\in(\mathbf{Q},\mathbf{K},\mathbf{V}),$ (3)
$\operatorname{SSA}^{\prime}(\mathbf{Q},\mathbf{K},\mathbf{V})=\mathrm{SN}\left(\mathbf{Q}\mathbf{K}^{\mathrm{T}}\mathbf{V}*s\right),$ (4)
where $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathcal{R}^{T\times N\times D}$, the spike-form $\mathbf{Q},\mathbf{K},\mathbf{V}$ are computed by learnable linear layers. $s$ is a scaling factor. $\mathrm{SN}$ means spiking neuron layer. The calculation of SSA avoids floating-point multiplication, meeting the property of SNNs.
@@ 3.2 Q-K Attention
An overview of Q-K attention is shown in Figure 1. Both VSA and SSA use three key components and have $O(N^{2}d)$ or $O(Nd^{2})$ computational complexity, while our proposed Q-K Attention which has linear complexity and only uses two spike-form components: $\mathbf{Q}$ and $\mathbf{K}$, which are produced through learnable linear matrics.
$\mathbf{Q}=\mathrm{SN}_{Q}\left(\mathrm{BN}\left(\mathbf{X}\mathbf{W}_{Q}\right)\right),\mathbf{K}=\mathrm{SN}_{K}\left(\mathrm{BN}\left(\mathbf{X}\mathbf{W}_{K}\right)\right),$ (5)
@@
where $\mathbf{X}$ is the input spiking map. According to the detailed calculation mechanism of $\mathbf{Q},\mathbf{K}$, Q-K Attention can be divided into Q-K Token Attention (QKTA) and Q-K Channel Attention (QKCA).
@@
Q-K Token Attention. We here assume $T=1$ and single head attention for mathematical description. After obtaining spike-form $\mathbf{Q},\mathbf{K}\in\mathcal{R}^{T\times N\times D}$, both $\mathbf{Q}$ and $\mathbf{K}$ can be formed as a spike matrix $N\times D$ ($N$ is the token number, $D$ is the channel number). QKTA process can be formulated as follows:
${\mathbf{A}}_{t}=\mathrm{SN}(\sum_{i=0}^{D}\mathbf{Q}_{i,j}),\quad\mathbf{X}^{\prime}={\mathbf{A}}_{t}\otimes\mathbf{K},$ (6)
@@
where ${\mathbf{A}}_{t}$ is the $N*1$ token attention vector, which models the binary importance of different tokens. ${\mathbf{A}}_{t}$ is a spike-form vector, which is obtained by addition operations (row summation) of $\mathbf{Q}$ spike matrix and a following spiking neuron. $\otimes$ is the Hadamard product between spike tensors, which is equivalent to the mask operation. We apply the spike-form token attention vector ${\mathbf{A}}_{t}$ to the $\mathbf{K}$ spike matrix through the column mask operation (token mask), to obtain the output $\mathbf{X}^{\prime}$ of QKTA.
@@
Q-K Channel Attention. The calculation process of Q-K channel attention is similar to the previous Q-K token attention, and can be formulated as : $\mathbf{A}_{c}=\mathrm{SN}(\sum_{j=0}^{N}\mathbf{Q}_{i,j}),\quad\mathbf{X}^{\prime}=\mathbf{A}_{c}\otimes\mathbf{K},$ where $\mathbf{A}_{c}$ is the $1*D$ channel attention vector, which models the binary importance of different channels. ${\mathbf{A}}_{t}$ is a spike-form vector, which is obtained by addition operations (column summation) of Q spike matrix and a following spiking neuron. Then, the output $\mathbf{X}^{\prime}$ of Q-K Channel Attention is obtained by the row mask operation (channel mask) between ${\mathbf{A}}_{t}$ and $\mathbf{K}$.
$\mathbf{X}^{\prime\prime}=\mathrm{SN}\left(\mathrm{BN}\left({\mathrm{Linear}}\left(\mathbf{X}^{\prime}\right)\right)\right).$ (7)
@@
As shown in Formula.7, a post-linear layer is also required after obtaining $\mathbf{X}^{\prime}$ of Q-K Token or Channel Attention. In addition, the channel dimension is $D/h$ in the multi-head Q-K attention, where $h$ is the head number. In this work, the spiking neuron uses the LIF model [28]. Same with [11], time step $T$ is an independent dimension for the spiking neuron layer. In other layers, it is merged with the batch size. We exploit QKTA in our experiments by default.
Linear Computational Complexity of Q-K Attention. As shown in Table 1, the time complexity of Q-K attention varies based on the implementation approach. Specifically, when utilizing spike-form broadcasted element-wise multiplication, $\otimes$, the time complexity can reach up to $O(N*D)$. When applying mask operation, the time complexity of Q-K attention is only $O(N)$ or $O(D)$. The space complexity of Q-K attention with the whole process is $O(N*D)$ at most, which is caused by the self-storage consumption Q and K matrix. In terms of the space complexity of attention operation, Q-K attention only requires an extra $1*D$ or $N*1$ space to store the attention vector with the space complexity of $O(N)$ or $O(D)$. The linear complexity of Q-K attention makes it possible to successfully explore the large-scale hierarchical architecture SNN model.
High Energy Efficiency of Q-K Attention. As a spike-driven attention module, the linear multiplication is transformed into sparse addition. Mask operation can be implemented on neuromorphic chips through addressing algorithms [29] or AND logic operations[30] with negligible power consumption. Compared with SSA, Q-K attention is much more energy-efficient, which comes from the following reasons: i) Q-K attention only adopts two spike-form components for spike [0, 1] operation without the V input and thus has less synaptic computing. ii) Q-K attention has much fewer spiking matrix operations due to its linear complexity of $O(N)$ or $O(D)$. iii) Q-K attention discards the scale operation of SSA, which leads to reduced power consumption further.
@@
Methods VSA [3] SSA [11] SDSA [13] QKTA QKCA
Time complexity $O(N^{2}D)$ $O(N^{2}D)$ $O(ND)$ $O(D)$ $O(N)$
Space complexity $O(N^{2}+ND)$ $O(N^{2}+ND)$ $O(ND)$ $O(N)$ $O(D)$
3.3 No Scaling Factors in Q-K Attention
In VSA [3], assume that $\mathbf{q}_{i}\left(\mathbf{q}_{i}\in R^{1\times d},\mathbf{Q}\in R^{m\times d}\right)$ and $\mathbf{k}_{i}\left(\mathbf{k}_{i}\in R^{1\times d},\mathbf{K}\in R^{m\times d}\right)$ are independent random variables with a mean of 0 and a variance of 1, then each element in the product of $\mathbf{Q}\mathbf{K}^{\mathrm{T}}$ has mean 0 and variance $d$. The variance magnitude of $\mathbf{Q}\mathbf{K}^{\mathrm{T}}$ grows with the embedding dimension $d$, which can result in gradient vanishing issues after softmax operation. Therefore, The product of matrices $\mathbf{Q}$ and $\mathbf{K}$ in VSA [3] is scaled by a factor $\frac{1}{\sqrt{d}}$ in Eq. 2 to normalize the product to variance 1. Though the softmax function is not adopted due to its non-spike operations (division, exponential operation) in SNNs, SSA-based [11] SNNs will suffer obvious performance degradation even cannot converge without scaling because the variance of $\mathbf{Q}\mathbf{K}^{\mathrm{T}}\mathbf{V}$ output is too large (Assuming that all the spiking elements are independent random variables and subject to Bernoulli Distribution). However, Q-K attention can discard scaling operations thus reducing power consumption because the variance of Q-K attention is much smaller than SSA (e.g. the max theoretical variance of Q-K token attention is only about 1 / 200 of SSA). The detailed analysis can be found in the Appendix.6.2 and Section.4.3.